# Can the Waymo Open Motion Dataset Support Realistic Behavioral Modeling? A Validation Study with Naturalistic Trajectories

### v1

@@ 3 Intersection Discharge Headway Analysis
Understanding the vehicles’ discharge headways at an intersection is crucial because it enables understanding of the road capacity. WOMD features numerous intersections where a large sample of starting-up trajectories of AVs and HVs are collected. Accordingly, WOMD is an appealing source to examine the effects of AVs on road capacity, which is a conclusion drawn by numerous existing studies through theoretical derivation and simulation [17, 18, 19]. Unfortunately, due to its data-smoothing technique, WOMD may fail to capture the discontinuity in the driving behavior of vehicles near intersections. Therefore, it is unclear whether WOMD enables unbiased analysis of vehicles’ discharging behavior when the signal turns green.
In this section, we assess the accuracy of through-movement discharge headways in WOMD by comparing the overall distribution to the PHX dataset and the headway compression phenomenon to an existing study [20].
@@ 3.1 Discharge Headway Extraction
To make sure that the collected headways are representative of the discharging process, we ensure that the collected time when vehicles start from intersections are from those that are already waiting at the intersection. This is to prevent biasing our analysis with free-flow headways. In addition, lanes that allow left-turns and the portion of queues behind right-turning vehicles are excluded, since turning vehicles tend to significantly lower their speeds and adopt longer headways [21]. We also note that due to the lack of signal information, we extract discharge headways starting from the second vehicle in the queue. Given the distinct data collection methods, headway extraction methods are customized for PHX and WOMD datasets.
@@ 3.1.1 PHX Dataset
In all, we identified 30 through-movement lanes with at least 2 discharging vehicles, and 10 of these lanes contain an AV. Figure 3 shows the geometry of an intersection in the PHX dataset. For each vehicle, we obtain the frame when its front bumper is closest to the red line in Figure 3. The stop line is not used as the reference line, because some vehicles’ front bumpers may be downstream of it while waiting. After obtaining the start times, we calculate the discharge headway by subtracting those of the adjacent vehicles.
@@
3.1.2 WOMD
In WOMD, we compare the smoothed trajectories to the map geometry file. First, we extract the through-movement lanes upstream of the intersections and identify vehicles that are on those at the first data collection time step. Since the center line of each lane is represented by densely spaced Cartesian coordinates, we map the location of these vehicles onto the closest centerline point and sort them based on their distances to the end of the through-movement lane. Then, we use whether the most downstream vehicle is initially stopped as an indicator to determine the signal state and discard all through-movement vehicles where the first vehicle is initially moving. Finally, we iterate over the initially-stopped vehicles and obtain the time when their front bumper passes through the stop-line. The vehicles’ front bumper locations can be obtained from the center locations, length, and direction of travel. If any of the following conditions are met, we discard that vehicle and upstream vehicles in the queue:
  1. 1.
    The vehicle never passes the stop line due to a data collection error, e.g., some vehicles may not be consistently identified in the dataset, and
  2. 2.
    The vehicle makes a turn after passing the stop line; this can be determined by extracting the next lane of the vehicle.
Based on the criteria above, we have extracted a total of 1152 lanes with discharging vehicles, where 180 lanes contain a Waymo vehicle.
@@ 3.2 Discharge Headway Distribution
With the collected headway data, we compare their distribution across three cases: HV following HV (HV-HV), AV following HV (AV-HV), and HV following AV (AV-HV). The sample size across the two datasets differs: In the PHX dataset, the sample sizes of (HV-HV), (AV-HV), and (AV-HV) are 110, 7, and 10, respectively. In WOMD, these numbers are 1788, 169, and 73, respectively. The overall headway distributions of the two datasets are shown in Figure 4.
To compare these distributions statistically, we perform the Kolmogorov-Smirnov test [22], which determines whether each case of discharge headway is drawn from the same headway distribution. The resulting p-values of HV-HV, AV-HV, and HV-AV cases are 0.125, 0.000, and 0.172, respectively, which means that under a 5% confidence level, we can only conclude that the headway distributions of the AV-HV case are different between the datasets. Indeed, according to Fig.4, AV-HV headway seems to be deterministic around 2s in the PHX dataset, which is lower than the lowest value observed in WOMD. This discrepancy indicates that WOMD might inaccurately capture AV trajectories or reflect outdated AV discharging behaviors, which are noticeably more conservative than those observed in the PHX dataset collected in the same year.
@@
While observing the overall headway distribution seems to suggest that the discharging behavior of HVs is similar between WOMD and PHX, examining the distribution of headway at different queue positions is important because the its mean tends to decrease for vehicles upstream of a queue [20, 21, 23]. This phenomenon is called headway compression and is observed in WOMD for HV-HV case, as shown by Fig.5.
Due to the small PHX sample size, we compare the HV-HV headways in WOMD to those in Lin et al. [20], where three intersections in the United States are included. Their reported results are reliable because they are obtained through traffic camera videos at a high resolution.
We conduct Welch’s t-tests to determine whether the population means of the headways are equal for the two datasets at each queue position [24]. Table 1 shows sample headway distribution and testing results. We note that the headway statistics for Lin et al.[20] are aggregated across all intersections.The p-values are less than 0.05 for all positions and the mean headways for WOMD are consistently higher. This indicates that WOMD may have overestimated the discharge headways for HV-HV case.
Position Lin et al. [20] WOMD t-score p-val
@@
Table 1: Headway Comparison between WOMD and Lin et al.[20] by Position in Queue
@@
4 Car-following Behavior Analysis
When comparing car-following behavior in PHX and WOMD, it is important to account for the high variability in human responses versus the relatively deterministic nature of AV’s control. Therefore, in this section, the AV car-following cases are extracted to perform the comparison. Recognizing that AV control policies may evolve across time and vary among locations, we use the March 2023 version of WOMD, which is the closest version to the PHX data collection period. We focus specifically on the decelerating-to-stop phase, which captures salient control actions such as gap perception, signal anticipation, and braking. This phase offers a sensitive lens for validating whether WOMD reflects real-world AV behavior.
@@ 4.1 Identification of Decelerating-to-Stop Cases
To identify car-following episodes from both the PHX and WOMD datasets, we applied consistent extraction criteria to ensure a fair comparison. Specifically, we extracted car-following episodes with a minimum duration of 10 seconds, taking into account the 20-second maximum trajectory length available in the WOMD dataset. We limited the maximum distance between vehicles during the car-following to 50 meters and excluded the car-following episodes where the following vehicle’s maximum speed during the episode did not exceed 3 $m/s$. This process yielded 143 car-following episodes from the PHX dataset and 47,826 pairs from the WOMD dataset. Among these, 16 pairs in PHX and 4,198 pairs in WOMD involved Waymo as the following vehicle.
Among the identified car following episodes, we focused on a specific subset in which the following vehicle decelerates to a stop behind the leader, extracting these segments using a specified process. The process begins by identifying time windows where both the follower and leader are moving slower than $\alpha$ for at least $\beta$ seconds. Once a stopping window for longer than $\beta$ seconds is found, we search backward in time to determine the onset of deceleration. We include 1 second from the onset of the stop of the following vehicle from this stopping window for each of the extracted segments of decelerating-to-stop cases. Each identified decelerating-to-stop segment includes up to $\gamma_{max}-1$ seconds of the follower vehicle’s deceleration leading up to the stop, plus 1 second of the follower vehicle being stopped. The segments shorter than $\gamma_{min}$ are discarded. To confirm that the follower’s stop results from the leading vehicle and not external factors like signals or pedestrians, segments where the spacing exceeds $\delta$ during the stop are also discarded. We set the thresholds as $\alpha=1.0m/s$, $\beta=1.0s$, $\gamma_{min}=3.0s$, $\gamma_{max}=10.0s$, and $\delta=4.0m$. After applying the process, we extracted 38 valid decelerating-to-stop segments from the PHX dataset, comprising 3,434 time steps (each 0.1 second), and 775 segments from the WOMD, totaling 48,909 time steps. Since the same extraction process was applied for both datasets, we expect the decelerating-to-stop trajectories of Waymo vehicles to align across the datasets. Figures 6(a–b) show example trajectories of the extracted decelerating-to-stop segments from the PHX and WOMD datasets.
@@ 4.2 Car-following Comparison Between PHX and WOMD using Markov Decision Process
To evaluate the behavior of Waymo vehicles in the PHX and WOMD datasets, we model the decelerating-to-stop dynamics as a Markov decision process. We treat the WOMD dataset as the ground truth, assuming it has no measurement error due to its high-fidelity data collection and large sample size. In contrast, we acknowledge measurement errors in the PHX dataset, which we address using the probabilistic methods described below. Given the extensive WOMD dataset, we train the transition matrix on WOMD and use it to assess whether PHX trajectories exhibit similar decelerating-to-stop behavior. This approach captures the probabilistic transitions between bins defined by relative speed ($\Delta v$), spacing ($g$), and follower speed ($v_{f}$). In the following, we detail the steps for constructing the transition matrix using WOMD, computing the transition probabilities for individual state transitions, and calculating the geometric mean probability per step for the entire decelerating-to-stop segment.
@@
where $\Phi_{\Delta v}$, $\Phi_{g}$, and $\Phi_{v_{f}}$ are the CDFs of the respective Gaussian error models for each variable.
@@ 4.3 Simulation–Extrapolation Adjustment of Car-Following Dynamics Comparison Using PHX Error Distribution
Trajectory data collected with onboard perception systems invariably contain measurement error. In this study, the PHX dataset comes with an empirically characterized error distribution, denoted $\widehat{F}_{\varepsilon}$, whereas the WOMD is treated in the literature as ground truth data [5, 25]. A formal bias correction procedure is required to determine whether the observed behavioral discrepancies between PHX and WOMD persist after the removal of PHX measurement noise. The SIMEX algorithm [13] fulfills this role: it artificially inflates the known error, estimates the bias in a chosen statistic, and then extrapolates that bias back to zero error. If, at the error-free point, the PHX–WOMD distance remains substantially larger than WOMD’s own internal variability, it suggests that the PHX dataset contains behavioral patterns absent from WOMD.
SIMEX was introduced by [13] to correct attenuation bias in linear‐regression coefficients measured with error. The method rapidly gained popularity in epidemiology, chemometrics, and econometrics, where instrument error is ubiquitous [26]. Empirical‐SIMEX, proposed by [27], replaces the parametric error generator with a non‐parametric bootstrap from the empirical error sample, allowing heavy tails and skewness. Subsequent work has applied SIMEX to generalized linear models and survival analysis [28], demonstrating its versatility wherever an explicit error distribution can be estimated.
@@
Due to the fact that car-following cases vary in duration and driver reaction lag, a frame-wise Euclidean distance comparison would confound behavioral differences with temporal shifts in drivers’ reactions. Dynamic Time Warping (DTW) resolves this by aligning two trajectories through a monotone warping path that minimizes cumulative local cost. While the classic formulation applies to scalar sequences [29, 30], we extend it to the multivariate state vector $\mathbf{y}(t)$ defined in Equation (6). Recent work shows that multivariate DTW captures driving behavior more robustly than single-variate measurements [31, 32].
@@
where $W=\operatorname{diag}\!\bigl{\{}\sigma_{g}^{-2},\sigma_{\Delta v}^{-2},\sigma_{v}^{-2}\bigr{\}}$ standardizes the three channels by their pooled variance. A warping path $W=[(i_{1},j_{1}),\dots,(i_{K},j_{K})]\subset[1{:}m]\times[1{:}n]$ must satisfy the boundary, continuity, and monotonicity conditions constraints of Sakoe–Chiba. The multivariate DTW distance is therefore
@@
which reduces to a shortest-path problem on the cumulative-cost matrix and is solvable in $O(mn)$ time [33].
Because the car-following episodes analyzed here have unequal length, We report the length-normalized DTW score $DTW^{\ast}=DTW/K,$ where $K$ is the path length, following the normalization proposed by [32]. This normalized distance enters the SIMEX procedure described above; by construction, it is invariant to constant reaction-time offsets and robust to differing episode durations.

### v2

@@ 3 Intersection Discharge Headway Analysis
Understanding the vehicles’ discharge headways at an intersection is crucial because it enables understanding of the road capacity. WOMD features numerous intersections where a large sample of starting-up trajectories of AVs and HVs are collected. Accordingly, WOMD is an appealing source to examine the effects of AVs on road capacity, which is a conclusion drawn by numerous existing studies through theoretical derivation and simulation (Yu et al., 2023; Park et al., 2021; Song and Fan, 2023). Unfortunately, due to its data-smoothing technique, WOMD may fail to capture the discontinuity in the driving behavior of vehicles near intersections. Therefore, it is unclear whether WOMD enables an unbiased analysis of vehicles’ discharging behavior when the signal turns green.
In this section, we assess the accuracy of through-movement discharge headways in WOMD by comparing the overall distribution to the PHX dataset and the headway compression phenomenon to an existing study (Lin and Thomas, 2005).
@@ 3.1 Discharge Headway Extraction
To make sure that the collected headways are representative of the discharging process, we ensure that the collected time when vehicles start from intersections is from those that are already waiting at the intersection. This is to prevent biasing our analysis with free-flow headways. In addition, lanes that allow left-turns and the portion of queues behind right-turning vehicles are excluded, since turning vehicles tend to significantly lower their speeds and adopt longer headways (Manual, 2000). We also note that due to the lack of signal information, we extract discharge headways starting from the second vehicle in the queue. Given the distinct data collection methods, headway extraction methods are customized for the PHX and WOMD datasets.
@@ 3.1.1 PHX Dataset
In all, we identified 30 through-movement lanes with at least two discharging vehicles, and 10 of these lanes contain an AV. Figure 3 shows the geometry of an intersection in the PHX dataset. For each vehicle, we obtain the frame when its front bumper is closest to the red line in Figure 3. The stop line is not used as the reference line, because some vehicles’ front bumpers may be downstream of it while waiting. After obtaining the start times, we calculate the discharge headway by subtracting those of the adjacent vehicles. This procedure yields headway measurements whose uncertainty is effectively bounded by the frame interval on the order of 0.1 s and the pixel resolution of the stabilized 8K imagery.
@@
3.1.2 Waymo Open Motion Dataset
In WOMD, we compare the smoothed trajectories to the map geometry file. First, we extract the through-movement lanes upstream of the intersections and identify vehicles that are on those at the first data collection time step. Since the center line of each lane is represented by densely spaced Cartesian coordinates, we map the location of these vehicles onto the closest centerline point and sort them based on their distances to the end of the through-movement lane. Then, we use whether the most downstream vehicle is initially stopped as an indicator to determine the signal state and discard all through-movement vehicles where the first vehicle is initially moving. Finally, we iterate over the initially-stopped vehicles and obtain the time when their front bumper passes through the stop-line. The vehicles’ front bumper locations can be obtained from the center locations, length, and direction of travel. If any of the following conditions are met, we discard that vehicle and upstream vehicles in the queue: (i) the vehicle never passes the stop line due to a data collection error, for example, some vehicles may not be consistently identified in the dataset, and (ii) the vehicle makes a turn after passing the stop line; this can be determined by extracting the next lane of the vehicle.
Based on the criteria above, we have extracted a total of 1,152 lanes with discharging vehicles, where 180 lanes contain a Waymo vehicle.
@@ 3.2 Discharge Headway Distribution
With the collected headway data, we compare their distribution across three cases: HV following HV (HV-HV), AV following HV (AV-HV), and HV following AV (AV-HV). The sample size across the two datasets differs: In the PHX dataset, the sample sizes of (HV-HV), (AV-HV), and (AV-HV) are 110, 7, and 10, respectively. In WOMD, these numbers are 1,788, 169, and 73, respectively. The overall headway distributions of the two datasets are shown in Figure 4.
To compare these distributions statistically, we perform the Kolmogorov-Smirnov test (Massey Jr, 1951), which determines whether each case of discharge headway is drawn from the same headway distribution. The resulting p-values of HV-HV, AV-HV, and HV-AV cases are 0.125, 0.000, and 0.172, respectively, which means that under a 5% confidence level, we can only conclude that the headway distributions of the AV-HV case are different between the datasets. Indeed, according to Fig.4, AV-HV headway seems to be deterministic around 2s in the PHX dataset, which is lower than the lowest value observed in WOMD. This discrepancy indicates that WOMD might inaccurately capture AV trajectories or reflect outdated AV discharging behaviors, which are noticeably more conservative than those observed in the PHX dataset collected in the same year.
@@
While observing the overall headway distribution seems to suggest that the discharging behavior of HVs is similar between WOMD and PHX, examining the distribution of headway at different queue positions is important because the its mean tends to decrease for vehicles upstream of a queue (Lin and Thomas, 2005; Manual, 2000; Greenshields et al., 1946). This phenomenon is called headway compression and is observed in WOMD for HV–HV case, as shown by Fig.5.
Due to the small PHX sample size, we compare the HV-HV headways in WOMD to those in Lin et al. (Lin and Thomas, 2005), where three intersections in the United States are included. Their reported results are reliable because they are obtained through traffic camera videos at a high resolution.
We conduct Welch’s t-tests to determine whether the population means of the headways are equal for the two datasets at each queue position (Welch, 1947). Table 1 shows the sample headway distribution and testing results. We note that the headway statistics for Lin et al.(Lin and Thomas, 2005) are aggregated across all intersections. The p-values are less than 0.05 for all positions, and the mean headways for WOMD are consistently higher. This indicates that WOMD may have overestimated the discharge headways for the HV-HV case.
Position Lin and Thomas (2005) WOMD t-score p-val
@@
Table 1: Headway Comparison between WOMD and Lin et al.(Lin and Thomas, 2005) by Position in Queue
@@
We note that intersection discharge dynamics inherently capture the acceleration regime of urban driving. The start-up process following a green signal corresponds to the acceleration-from-stop phase, complementing the deceleration-to-stop analysis presented later in Section 4.
4 Deceleration-to-Stop Behavior Analysis
When comparing car-following behavior in PHX and WOMD, it is important to account for the high variability in human responses versus the relatively deterministic nature of AV’s control. Therefore, in this section, we focus on the decelerating-to-stop phase of car-following, which captures salient control actions such as gap perception, signal anticipation, and braking. This phase offers a sensitive lens for validating whether WOMD reflects real-world AV behavior.
@@ 4.1 Identification of Decelerating-to-Stop Cases
To identify car-following episodes from both the PHX and WOMD datasets, we applied consistent extraction criteria to ensure a fair comparison. Specifically, we extracted car-following episodes with a minimum duration of 10 seconds, taking into account the 20-second maximum trajectory length available in the WOMD dataset. We limited the maximum distance between vehicles during the car-following to 50 meters and excluded the car-following episodes where the following vehicle’s maximum speed during the episode did not exceed 3 $m/s$. This process yielded 143 car-following episodes from the PHX dataset and 47,826 pairs from the WOMD dataset. Among these, the WOMD scenarios contain 4,198 episodes where the following vehicle is flagged as the self-driving car (SDC) in WOMD metadata, whereas all ego vehicles in the PHX dataset correspond to the driverless ride-hail service described in Section 2.1.
Among the identified car following episodes, we focused on a specific subset in which the following vehicle decelerates to a stop behind the leader, extracting these segments using a specified process. The process begins by identifying time windows where both the follower and leader are moving slower than $\alpha$ for at least $\beta$ seconds. Once a stopping window for longer than $\beta$ seconds is found, we search backward in time to determine the onset of deceleration. We include 1 second from the onset of the stop of the following vehicle from this stopping window for each of the extracted segments of decelerating-to-stop cases. Each identified decelerating-to-stop segment includes up to $\gamma_{max}-1$ seconds of the follower vehicle’s deceleration leading up to the stop, plus 1 second of the follower vehicle being stopped. The segments shorter than $\gamma_{min}$ are discarded. To confirm that the follower’s stop results from the leading vehicle and not from external factors such as signals or pedestrians, segments in which the spacing exceeds $\delta$ during the stop are also discarded. We set the thresholds as $\alpha=1.0m/s$, $\beta=1.0s$, $\gamma_{min}=3.0s$, $\gamma_{max}=10.0s$, and $\delta=4.0m$. After applying the process, we extracted 38 valid decelerating-to-stop segments from the PHX dataset, comprising 3,434 time steps (each 0.1 second), and 775 segments from the WOMD, totaling 48,909 time steps. Since the same extraction process was applied for both datasets, we expect the decelerating-to-stop trajectories of Waymo vehicles to align across the datasets. Figures 6(a–b) show example trajectories of the extracted decelerating-to-stop segments from the PHX and WOMD datasets.
@@
To avoid artifacts from occasional tracking failures in both datasets, we applied an outlier filter before the Markov and DTW analyses. Any car-following episode containing a time step with longitudinal acceleration magnitude exceeding $10~\mathrm{m/s^{2}}$ or an instantaneous change in spacing greater than $30~\mathrm{m}$ over a single $0.1$ s interval was discarded. This filter removed fewer than $1\%$ of candidate episodes in each dataset and did not significantly affect the summary statistics reported below.
@@ 4.2 Car-following Comparison Between PHX and WOMD using Markov Decision Process
To evaluate the behavior of Waymo vehicles in the PHX and WOMD datasets, we model the decelerating-to-stop dynamics as a Markov decision process. We treat the WOMD dataset as a high-fidelity behavioral reference, in the sense that it offers a large number of episodes collected with on-board sensing and proprietary tracking, but we do not assume that WOMD trajectories are entirely free of smoothing or perception error. In contrast, the PHX dataset exhibits empirically characterized error due to aerial videography and trajectory extraction, which we explicitly account for using the probabilistic methods described below. Given the extensive WOMD dataset, we train the transition matrix on WOMD and use it to assess whether PHX trajectories exhibit similar decelerating-to-stop behavior.
@@
where $\Phi_{\Delta v}$, $\Phi_{g}$, and $\Phi_{v_{f}}$ are the CDFs of the respective Gaussian error models for each variable. The above construction assumes that the three state variables are independent given the observation, so that the joint probability over a bin factorizes into the product of marginal probabilities. This independence approximation is reasonable in our setting because the empirically estimated correlations in the PHX error model are modest. A more general formulation would incorporate the full empirical covariance matrix when computing bin probabilities in $q_{t}$, which we leave as an extension for future work.
@@ 4.3 Simulation–Extrapolation Adjustment of Car-Following Dynamics Comparison Using PHX Error Distribution
Trajectory data collected with onboard perception systems invariably contain measurement error. In this study, the PHX dataset comes with an empirically characterized error distribution, denoted $\widehat{F}_{\varepsilon}$, whereas WOMD is often used as if it were ground-truth data in the literature(Hu et al., 2022, 2023b). In reality, WOMD trajectories are themselves the output of an on-board perception and temporal-smoothing pipeline, for which no explicit error characterization is available. This asymmetry implies that we can only apply SIMEX to the PHX dataset, for which an empirical error sample has been obtained, and must treat any latent measurement error in WOMD as part of the reference distribution against which we compare.
A formal bias correction procedure is required to determine whether the observed behavioral discrepancies between PHX and WOMD persist after the removal of PHX measurement noise. The SIMEX algorithm (Cook and Stefanski, 1994) fulfills this role: it artificially inflates the known error, estimates the bias in a chosen statistic, and then extrapolates that bias back to zero error. If, at the error-free point, the PHX–WOMD distance remains substantially larger than WOMD’s own internal variability, it suggests that the PHX dataset contains behavioral patterns absent from WOMD.
SIMEX was introduced by Cook and Stefanski (1994) to correct attenuation bias in linear‐regression coefficients measured with error. The method rapidly gained popularity in epidemiology, chemometrics, and econometrics, where instrument error is ubiquitous (Carroll et al., 1995). Empirical‐SIMEX, proposed by Devanarayan and Stefanski (2002), replaces the parametric error generator with a non‐parametric bootstrap from the empirical error sample, allowing heavy tails and skewness. Subsequent work has applied SIMEX to generalized linear models and survival analysis (Hardin et al., 2003), demonstrating its versatility wherever an explicit error distribution can be estimated.
@@
Due to the fact that car-following cases vary in duration and driver reaction lag, a frame-wise Euclidean distance comparison would confound behavioral differences with temporal shifts in drivers’ reactions. Dynamic Time Warping (DTW) resolves this by aligning two trajectories through a monotone warping path that minimizes cumulative local cost. While the classic formulation applies to scalar sequences (Sakoe and Chiba, 1978; Myers et al., 1980), we extend it to the multivariate state vector $\mathbf{y}(t)$ defined in Equation (6). Recent work shows that multivariate DTW captures driving behavior more robustly than single-variate measurements (Hosseini et al., 2022; Zhang et al., 2025).
@@
where $W=\operatorname{diag}\!\bigl\{\sigma_{g}^{-2},\sigma_{\Delta v}^{-2},\sigma_{v}^{-2}\bigr\}$ standardizes the three channels by their pooled variance. The diagonal weighting matrix $W$ standardizes each channel by its pooled variance so that spacing, relative speed, and speed enter the local cost in comparable, dimensionless units. In principle, one could replace $W$ by the inverse of the full empirical covariance matrix, yielding a Mahalanobis-type distance that accounts for cross-channel correlations (MAHALANOBIS, 1936). We experimented with this alternative and found that it produces very similar rankings of DTW distances and, in particular, does not change the qualitative separation between PHX–WOMD and WOMD–WOMD pairs. Because the diagonal form is easier to interpret and less sensitive to covariance estimation error in small samples, we adopt it as the primary specification.
A warping path $W=[(i_{1},j_{1}),\dots,(i_{K},j_{K})]\subset[1{:}m]\times[1{:}n]$ must satisfy the boundary, continuity, and monotonicity conditions constraints of Sakoe–Chiba (Sakoe and Chiba, 1978). The multivariate DTW distance is therefore
@@
which reduces to a shortest-path problem on the cumulative-cost matrix and is solvable in $O(mn)$ time (Senin, 2008).
Because the car-following episodes analyzed here have unequal length, We report the length-normalized DTW score $DTW^{\ast}=DTW/K,$ where $K$ is the path length, following the normalization proposed by (Zhang et al., 2025). This normalized distance enters the SIMEX procedure described above; by construction, it is invariant to constant reaction-time offsets and robust to differing episode durations.